A small, spring-loaded cannon launches a golf ball from level ground with an initial speed vi at an angle θi with the horizontal. The golf ball lands a horizontal distance R from its launch point. The highest point the golf ball reaches during its flight is a distance R/12 above the ground. In terms of R and g, find the following. (You may ignore air resistance.)

(a)

the time interval during which the golf ball is in motion

​​​​​t =

(b)

the golf ball's speed at the peak of its path

v =

(c)

the initial vertical component of its velocity

vyi =

ONLY 3 PARTS FOR CREDIT. REST FOR PRACTICE.

(d)

its initial speed

(e)

the angle θi (Assume θi is in radians. Do not include units in your answer.)

(f)

Suppose the golf ball is launched at the same initial speed found in (d) but at the angle appropriate for reaching the greatest height that it can. Find this height.

(g)

Suppose the golf ball is launched at the same initial speed but at the angle for greatest possible range. Find this maximum horizontal range.
Answer #1

Given given, initial relouty of her projection = n Angle of projection, Op. Range =R, maximum height, Hmax e (a) Time Entequawe know, range, R = u Sin 20; - Sus 20; 9 g = v = Rg Rg = Ve. Sin 20. 2 Sino costi T= 2u sunde = a (ne sander) but R Tano: =(c) gritual vertical component of relocity is Vyi = Ve Sinoo & Sunde / RgT V asiso, coso; voi = Rg tandi = Rg (3) Vaji = v Ro(f) for maximum height, 0290 Then Maximum height reached is. Hmax = & sing 29 5Rg Hmax = 58 (9) for Range to be maximum, og =